OpenStudy (anonymous):

Which is the graph of the line with equation: y – 4 = 2(x + 1)

OpenStudy (anonymous):

Please post graphs.

OpenStudy (anonymous):

I would highly recommend getting your equation in y = mx + b form by expanding any expressions that can be expanded and getting y on a side by itself.

OpenStudy (anonymous):

So! Put the equation into y=mx+b form! It really, really helps bring clarity to what you are looking for.

OpenStudy (anonymous):

Do you need help accomplishing this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So first off, and easiest, let's get y on a side by itself.

OpenStudy (anonymous):

y – 4 = 2(x + 1) Can you please get rid of that -4 so that y is on its own side?

OpenStudy (anonymous):

Use algebra.